# Integrating Sequential and Relational Modeling for User Events: Datasets and Prediction Tasks

### v1

@@ Abstract
User event modeling plays a central role in many machine learning applications, with use cases spanning e-commerce, social media, finance, cybersecurity, and other domains. User events can be broadly categorized into personal events, which involve individual actions, and relational events, which involve interactions between two users. These two types of events are typically modeled separately, using sequence-based methods for personal events and graph-based methods for relational events. Despite the need to capture both event types in real-world systems, prior work has rarely considered them together. This is often due to the convenient simplification that user behavior can be adequately represented by a single formalization, either as a sequence or a graph. To address this gap, there is a need for public datasets and prediction tasks that explicitly incorporate both personal and relational events. In this work, we introduce a collection of such datasets, propose a unified formalization, and empirically show that models benefit from incorporating both event types. Our results also indicate that current methods leave notable room for improvements. We release these resources [^1] Link to the datasets and prediction tasks. Link to the data construction and experimentation code. to support further research in unified user event modeling and encourage progress in this direction.
@@ 1 Introduction
Modeling user events is a central task in machine learning with broad applications across various domains [1, 2]. In e-commerce, it is used to capture user preferences for personalized ranking and product recommendation [3]. In social media platforms, event modeling supports feed optimization and engagement prediction by inferring user interests over time [4]. Financial systems leverage user behavior data for fraud detection, credit risk assessment, and behavioral profiling [5]. Online services such as search and streaming platforms rely on user event sequences for content recommendation under real-time constraints [6, 7]. In cybersecurity, modeling user and system events is essential for detecting anomalies and preventing intrusions [8, 9]. These applications demonstrate the importance of building models that can effectively capture complex, context-dependent user behavior from event sequences.
User events can be broadly categorized into personal and relational events. Personal events involve only a single user and reflect individual actions, such as searching for content, viewing items, or posting updates. In contrast, relational events involve interactions between two or more users, such as following another user, co-editing a document, or exchanging messages. Traditionally, these two types of events are often modeled separately. Relational events are commonly modeled using graph-based approaches that capture structural dependencies and interaction patterns among users [10, 11]. On the other hand, personal events are typically modeled as sequences using recurrent or attention-based architectures to capture temporal dependencies in personal event histories [12, 13, 14].
@@
There have been efforts in the graph area to capture both structural and temporal dependencies using temporal graph formalizations (such as CTDG [15, 16, 17]) and models built on top of these formalizations (such as TGAT [18], TGN [16], DyRep [19], TNCN [20], and others [21, 22, 23]). However, these approaches primarily focus on the temporal dependencies of relational events while neglecting personal events. For example, the formalization used in the Temporal Graph Benchmark (TGB) papers [24, 25] and recent temporal graph models [20, 21, 22, 23] defines a temporal graph as a stream of triplets consisting of source, destination, and timestamp. Personal events that involve only a single entity cannot be directly represented under this formulation. One workaround is to convert all personal events into nodes and define personal events as triplets of user node, event node, and timestamp. However, this construction is not as straightforward for capturing temporal dependencies in personal event histories compared to sequence-based modeling.
@@
  • We invite the research community to use these resources and help close the gap in unified user event modeling.
@@ 2 Related Works
Event sequence modeling, such as temporal point processes [26, 27, 28], deals with event stream data $(e_{1},t_{1}),(e_{2},t_{2}),\cdots$, where each $e_{i}$ is an event type drawn from an event set $\mathcal{E}=\{1,2,\cdots,|\mathcal{E}|\}$, and $0\leq t_{1}\leq t_{2}\leq\cdots$ are the times of occurrence. In user event sequence modeling, such as personalized event prediction tasks [29, 30, 31], we have $N$ users, denoted as $\mathcal{U}=\{u_{1},u_{2},\cdots,u_{N}\}$. Each user has their own sequence of events, denoted $\text{Seq}(u)=[(e_{1},t_{1})^{(u)},(e_{2},t_{2})^{(u)},\cdots]$. In some applications, exact timestamps are replaced with discrete time steps, simplifying the definition to $\text{Seq}(u)=[e_{1}^{(u)},e_{2}^{(u)},\cdots]$. In other variants of sequence modeling, the event space may be restricted to a single domain, for example, products consumed by a user in sequential recommendation tasks [32, 33, 13]. While these user event sequence models capture the complexity of event sequences within a user, they lack the ability to encode user-to-user interactions.
Graph modeling, on the other hand, explicitly encodes user-to-user interactions through its node and edge abstraction, $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ and $\mathcal{E}$ denote the node and edge sets, respectively. Temporal graph abstractions, such as CTDG [15, 16], incorporate temporal dynamics by representing a graph as a sequence of time-stamped events $\mathcal{G}=\{x(t_{1}),x(t_{2}),\cdots\}$. Each event $x(t)$ can be either (1) a node-wise event (e.g., node addition or feature update) represented by a feature vector $\mathbf{\bm{v}}_{i}(t)$, where $i$ is the node index, or (2) an interaction event between node $i$ and node $j$ represented by a temporal edge $\mathbf{\bm{e}}_{i,j}(t)$. Another CTDG formulation by Kazemi et al. [17] describes a temporal graph as a pair consisting of an initial graph and an observation set $(\mathcal{G},\mathcal{O})$. The observation set describes the evolution of the graph, with each observation represented as a tuple (event type, event, timestamp), where the event type may include edge addition, edge deletion, node addition, node deletion, node splitting, node merging, and so on. In practice, however, many temporal graph models [20, 21, 22, 23] omit node-level events altogether and instead define a temporal graph as a sequence of interaction events expressed as triplets, $\mathcal{G}=\{(u_{1},v_{1},t_{1}),(u_{2},v_{2},t_{2}),\cdots\}$, where $u$ and $v$ represent the source and destination nodes.
@@ 3 Problem Formalization
For our Personal and Relational User Event Sequence (PRES) modeling, we take inspiration from the standard user event sequence modeling. We denote the set of users (which can also be a customer, account, entity, etc.) as $\mathcal{U}=\{u_{1},u_{2},\cdots,u_{N}\}$, where $N$ is the number of users. Each user has their own sequence of events that occur over time. For example, the sequence for user $u_{i}$ is denoted as $\text{Seq}(u_{i})=[(e_{1},t_{1}),(e_{2},t_{2}),\cdots,(e_{M_{i}},t_{M_{i}})]$, where $e$ describes an event, $t$ describes the time at which the event occurs, and $M_{i}$ denotes the number of events for user $u_{i}$. Each user may have a different number of events in their event sequence. We denote the set of all user sequences by ${\mathcal{S}}=\left\{\text{Seq}(u)\mid u\in\mathcal{U}\right\}$. An event may come from two different event sets: the personal event set and the relational event set. The personal event set contains a set of events that can occur for an individual user; $p\in\mathcal{P}\triangleq\{1,2,\cdots,|\mathcal{P}|\}$. The relational event set contains a set of all possible events $r\in\mathcal{R}\triangleq\{1,2,\cdots,|\mathcal{R}|\}$, which involve a relation from one user to another. Thus, an event can be defined by a personal event $e=p$, or a relational event tuple $e=(r,v)$, where $v$ is another user.
@@
  • In the temporal graph formulation, events are ordered globally across all users, similar to standard (non-user-based) event sequence modeling. In contrast, PRES follows personalized user-event sequence modeling, where each user has their own separate sequence of events.
  • The temporal graph formulation places greater emphasis on modeling interaction events between two nodes, often neglecting user events that do not involve another user. Many recent temporal graph models, such as TNCN [20], NAT [21], DyGFormer [22], and others [23, 34, 35]; recent temporal graph benchmarks, such as TGB [24] and TGB 2.0 [25]; as well as popular graph frameworks such as PyTorch Geometric [36], even reduce the formulation to a sequence of triplets representing interaction events with no support for encoding personal events. By contrast, PRES treats personal events as first-class components. This is particularly relevant in many areas mentioned in the introduction, where the number of personal events is far larger than that of relational events.
@@
When we compare with the CTDG formulation by Kazemi et al. [17], all the concerns above remain. While this formulation allows more flexible observations, it is still restricted to events that describe the evolution of the graph through changes in nodes or edges, such as edge additions, edge deletions, node additions, node deletions, node splitting, or node merging.
@@ 4.1 Dataset Information
We curated user event datasets from multiple domains [^2] The datasets presented in this paper are used solely to demonstrate the claims made in this work and are not reflective of any datasets leveraged by Capital One., following recent graph dataset curation works [24, 25, 37, 38, 39], and then processed them according to our formalization in Section 3. The data is stored in CSV format with the columns: uid, timestamp, event_set, event, and other_uid (See Appendix B for details). The uid is a numerical user ID, whereas event_set indicates whether the event is personal or relational. For relational events, other_uid refers to the other user involved in the relation; for personal events, this column is null. The statistics of each datasets is shown in Table 1. More details on collection and dataset license are available in Appendix A. The data processing code are open-sourced and descriptions are available in Appendix C.
@@
pres-brightkite. This dataset contains location check-ins and friendship history of Brightkite users, a location-based social networking platform. It was originally collected by Cho et al. [40] and published in the SNAP Dataset Repository [41]. Personal events consist of sequences of location check-ins. We convert the original latitude and longitude coordinates into Geohash-8 representations [42, 43], short alphanumeric strings encoding geographic locations. Nearby locations share similar geohash prefixes, while distant ones differ. Example geohashes include 9v6kpmr1, gcpwkeq6, and u0yhxgm1. Relational events capture friendship connections among users. The dataset includes 58,228 users and 5,130,866 events. Only personal events have timestamps; relational events do not.
@@
pres-gowalla. The dataset also contains the location check-in and friendship history of another social network platform, Gowalla. It was also originally collected by Cho et al. [40] and published in the SNAP Repository [41]. We processed and formatted the data following the same approach used for pres-brightkite. The dataset contains personal events from geohash check-ins and relational events from friendship connections, totaling 8,342,943 events from 196,591 users.
@@
pres-amazon-clothing. The dataset contains Amazon product reviews and ratings in the Clothing, Shoes and Jewelry category, spanning from May 1996 to July 2014. The raw data was originally collected by McAuley et al. [44]. In this dataset, we define personal events as sequences of product IDs and ratings reviewed by a user, for example: B [^0] MLDCZ2:5 and B001OE3F08:3. Relational events represent co-review patterns, where two users have reviewed at least three of the same products. The dataset contains event sequences from 185,986 users, with a total of 1,591,947 events.
@@
pres-amazon-electronics. The dataset contains Amazon product reviews and ratings in the Electronics category, originally collected by McAuley et al. [44]. As in pres-amazon-clothing, personal events are defined as sequences of product IDs and ratings, while relational events capture co-review patterns. In total, the dataset contains 2,938,178 events from 254,064 users.
@@
pres-github. This dataset contains GitHub user activity from January 2025, collected from the GH Archive. Personal events include actions such as Push, CreateBranch, CreateRepository, PullRequestOpened, IssuesOpened, and Fork. Relational events represent project collaboration, where two users are linked if both contributed at least five commits or pull requests to the same repository. The dataset includes 102,878,895 events from 3,669,079 users. Only personal events include timestamps; relational events do not, similar to pres-brightkite and pres-gowalla.
@@
Variability of the datasets. As shown in Table 1, the PRES datasets vary significantly across multiple aspects. The number of users ranges from around 58 thousand in pres-brightkite to more than 3.5 million in pres-github. The number of events also varies, from approximately 1.5 million in pres-brightkite to over 100 million in pres-github. The ratio between relational and personal events ranges from around 1:3 in pres-gowalla to approximately 1:80 in pres-amazon-clothing. The number of unique events also differs widely, from just 24 in pres-github to more than 1 million in pres-gowalla. In addition, we observe variability in the number of users having personal events, relational events, and both. Some datasets have more users with relational events than with personal events (e.g., pres-brightkite, pres-gowalla), while others show the opposite trend (e.g., pres-amazon-clothing, pres-amazon-electronics, pres-github). These differences in dataset properties present distinct challenges for modeling user events in each dataset.

### v2

@@ Abstract
User event modeling plays a central role in many machine learning applications, with use cases spanning e-commerce, social media, finance, cybersecurity, and other domains. User events can be broadly categorized into personal events, which involve individual actions, and relational events, which involve interactions between two users. These two types of events are typically modeled separately, using sequence-based methods for personal events and graph-based methods for relational events. Despite the need to capture both event types in real-world systems, prior work has rarely considered them together. This is often due to the convenient simplification that user behavior can be adequately represented by a single formalization, either as a sequence or a graph. To address this gap, there is a need for public datasets and prediction tasks that explicitly incorporate both personal and relational events. In this work, we introduce a collection of such datasets, propose a unified formalization, and empirically show that models benefit from incorporating both event types. Our results also indicate that current methods leave notable room for improvements. We release these resources [^1] Link to the datasets and prediction tasks. Link to the dataset construction and experimentation code. to support further research in unified user event modeling and encourage progress in this direction.
@@ 1 Introduction
Modeling user events is a central task in machine learning with broad applications across various domains [purificato2024user, martin2021survey, deng2024advances]. In e-commerce, it is used to capture user preferences for personalized ranking and product recommendation [he2023survey, zeng2019user]. In social media platforms, event modeling supports feed optimization and engagement prediction by inferring user interests over time [abdelhafez2013survey, piao2018inferring, fang2023snapchat]. Financial systems leverage user behavior data for fraud detection, credit risk assessment, and behavioral profiling [hernandez2024fraud, mashrur2020risk, ozbayoglu2020deep, hojaji2022behavioral]. Online services such as search and streaming platforms rely on user event sequences for content recommendation under real-time constraints [zhao2023mbsrs, gayoavello2009session, covington2016youtube, amatriain2013netflix]. In cybersecurity, modeling user and system events is essential for detecting anomalies and preventing intrusions [lashkari2017survey, wang2017clickstream]. These applications demonstrate the importance of building models that can effectively capture complex, context-dependent user behavior from event sequences.
User events can be broadly categorized into personal and relational events. Personal events involve only a single user and reflect individual actions, such as searching for content, viewing items, or posting updates. In contrast, relational events involve interactions between two or more users, such as following another user, co-editing a document, or exchanging messages. Traditionally, these two types of events are often modeled separately. Relational events are commonly modeled using graph-based approaches that capture structural dependencies and interaction patterns among users [gao2023survey, guo2020survey, wu2022graph, li2024graph]. On the other hand, personal events are typically modeled as sequences using recurrent or attention-based architectures to capture temporal dependencies in personal event histories [Chen2024ASO, pan2024survey, boka2024survey, sun2019bert4rec, Tan2016ImprovedRN, Kang2018SelfAttentiveSR, liguori2023modeling].
@@
There have been efforts in the graph area to capture both structural and temporal dependencies using temporal graph formalizations (such as CTDG [nguyen2018continuous, tgn_icml_grl2020, kazemi2020representation]) and models built on top of these formalizations (such as TGAT [xu2020inductive], TGN [tgn_icml_grl2020], DyRep [trivedi2019dyrep], TNCN [zhang2406efficient], and others [luo2022neighborhood, yu2023towards, gravina2024long]). However, these approaches primarily focus on the temporal dependencies of relational events while neglecting personal events. For example, the formalization used in the Temporal Graph Benchmark (TGB) papers [huang2023temporal, gastinger2024tgb] and recent temporal graph models [zhang2406efficient, luo2022neighborhood, yu2023towards, gravina2024long] defines a temporal graph as a stream of triplets consisting of source, destination, and timestamp. Personal events that involve only a single entity cannot be directly represented under this formulation. One workaround is to convert all personal events into nodes and define personal events as triplets of user node, event node, and timestamp. However, this construction is not as straightforward for capturing temporal dependencies in personal event histories compared to sequence-based modeling.
@@
  • We invite the community to use these resources and close the gap in unified user event modeling.
@@ 2 Related Works
Event sequence modeling, such as temporal point processes [mei2017neural, zuo2020transformer, zhang2020self], deals with event stream data $(e_{1},t_{1}),(e_{2},t_{2}),\cdots$, where each $e_{i}$ is an event type drawn from an event set $\mathcal{E}=\{1,2,\cdots,|\mathcal{E}|\}$, and $0\leq t_{1}\leq t_{2}\leq\cdots$ are the times of occurrence. In user event sequence modeling, such as personalized event prediction tasks [letham2013sequential, boyd2020user, yang2022towards], we have $N$ users, denoted as $\mathcal{U}=\{u_{1},u_{2},\cdots,u_{N}\}$. Each user has their own sequence of events, denoted $\text{Seq}(u)=[(e_{1},t_{1})^{(u)},(e_{2},t_{2})^{(u)},\cdots]$. In some applications, exact timestamps are replaced with discrete time steps, simplifying the definition to $\text{Seq}(u)=[e_{1}^{(u)},e_{2}^{(u)},\cdots]$. In other variants of sequence modeling, the event space may be restricted to a single domain, for example, products consumed by a user in sequential recommendation tasks [hidasi2015session, kang2018self, sun2019bert4rec]. While these user event sequence models capture the complexity of event sequences within a user, they lack the ability to encode user-to-user interactions.
Graph modeling, on the other hand, explicitly encodes user-to-user interactions through its node and edge abstraction, $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ and $\mathcal{E}$ denote the node and edge sets, respectively. Temporal graph abstractions, such as CTDG [nguyen2018continuous, tgn_icml_grl2020], incorporate temporal dynamics by representing a graph as a sequence of time-stamped events $\mathcal{G}=\{x(t_{1}),x(t_{2}),\cdots\}$. Each event $x(t)$ can be either (1) a node-wise event (e.g., node addition or feature update) represented by a feature vector $\mathbf{\bm{v}}_{i}(t)$, where $i$ is the node index, or (2) an interaction event between node $i$ and node $j$ represented by a temporal edge $\mathbf{\bm{e}}_{i,j}(t)$. Another CTDG formulation by kazemi2020representation describes a temporal graph as a pair consisting of an initial graph and an observation set $(\mathcal{G},\mathcal{O})$. The observation set describes the evolution of the graph, with each observation is represented as a tuple (event type, event, timestamp), where the event type may include edge addition, edge deletion, node addition, node deletion, node splitting, node merging, etc. In practice, however, many temporal graph models [zhang2406efficient, luo2022neighborhood, yu2023towards, gravina2024long] omit node-level events altogether and instead define a temporal graph as a sequence of interaction events expressed as triplets, $\mathcal{G}=\{(u_{1},v_{1},t_{1}),(u_{2},v_{2},t_{2}),\cdots\}$, where $u$ and $v$ represent the source and destination nodes.
@@ 3 Problem Formalization
For our Personal and Relational User Event Sequence (PRES) modeling, we take inspiration from the standard user event sequence modeling. We denote a set of users (which can also be a customer, account, entity, etc.) as $\mathcal{U}=\{u_{1},u_{2},\cdots,u_{N}\}$, where $N$ is the number of users. Each user has their own sequence of events that occur over time. For example, the sequence for user $u_{i}$ is denoted as $\text{Seq}(u_{i})=[(e_{1},t_{1}),(e_{2},t_{2}),\cdots,(e_{M_{i}},t_{M_{i}})]$, where $e$ describes an event, $t$ describes the time at which the event occurs, and $M_{i}$ denotes the number of events for user $u_{i}$. Each user may have a different number of events in their event sequence. We denote the set of all user sequences by ${\mathcal{S}}=\left\{\text{Seq}(u)\mid u\in\mathcal{U}\right\}$. An event may come from two different event sets: the personal event set and the relational event set. The personal event set contains a set of events that can occur for an individual user; $p\in\mathcal{P}\triangleq\{1,2,\cdots,|\mathcal{P}|\}$. The relational event set contains a set of all possible events $r\in\mathcal{R}\triangleq\{1,2,\cdots,|\mathcal{R}|\}$, which involve a relation from one user to another. Thus, an event can be defined by a personal event $e=p$, or a relational event tuple $e=(r,v)$, where $v$ is another user.
@@
  • In the temporal graph formulation, events are ordered globally across all users, similar to the standard (non-user-based) event sequence modeling. In contrast, PRES follows personalized user-event sequence modeling, where each user has their own separate sequence of events.
  • The temporal graph formulation places greater emphasis on modeling interaction events between two nodes, often neglecting user events that do not involve another user. Many recent temporal graph models, such as TNCN [zhang2406efficient], NAT [luo2022neighborhood], DyGFormer [yu2023towards], and others [gravina2024long, lu2024improving, gao2025hyperevent]; recent temporal graph benchmarks, such as TGB [huang2023temporal] and TGB 2.0 [gastinger2024tgb]; as well as popular graph frameworks such as PyTorch Geometric [fey2019fast], even reduce the formulation to a sequence of triplets representing interaction events with no support for encoding personal events. By contrast, PRES treats personal events as first-class components. This is particularly relevant in many areas mentioned in the introduction, where the number of personal events is far larger than that of relational events.
@@
When we compare with the CTDG formulation by kazemi2020representation, all the concerns above remain. While this formulation allows more flexible observations, it is still restricted to events that describe the evolution of the graph through changes in nodes or edges, such as edge additions, edge deletions, node additions, node deletions, node splitting, or node merging.
@@ 4.1 Dataset Information
We curated user event datasets from multiple domains [^2] The datasets presented in this paper are used solely to demonstrate the claims made in this work and are not reflective of any datasets leveraged by Capital One., following recent graph dataset curation works [huang2023temporal, gastinger2024tgb, yi2025tgb, hu2020open, hu2021ogb], and then processed them according to our formalization in Section 3. The data is stored in CSV format with the columns: uid, timestamp, event_set, event, and other_uid (See Appendix B for details). The uid is a numerical user ID, whereas event_set indicates whether the event is personal or relational. For relational events, other_uid refers to the other user involved in the relation; for personal events, this column is null. The statistics of each datasets is shown in Table 1. More details on collection and dataset license are available in Appendix A. The data processing code are open-sourced and descriptions are available in Appendix C.
@@
pres-brightkite. This dataset contains location check-ins and friendship history of Brightkite users, a location-based social networking platform. It was originally collected by cho2011friendship and published in the SNAP Dataset Repository [snapnets]. Personal events consist of sequences of location check-ins. We convert the original latitude and longitude coordinates into Geohash-8 representations [niemeyer2008geohash, morton1966computer], short alphanumeric strings encoding geographic locations. Nearby locations share similar geohash prefixes, while distant ones differ. Example geohashes include 9v6kpmr1, gcpwkeq6, and u0yhxgm1. Relational events capture friendship connections among users. The dataset includes 58,228 users and 5,130,866 events. In this dataset, only the personal events include timestamp information; relational events do not have associated timestamps.
@@
pres-gowalla. The dataset also contains the location check-in and friendship history of another social network platform, Gowalla. It was also originally collected by cho2011friendship and published in the SNAP Repository [snapnets]. We processed and formatted the data following the same approach used for pres-brightkite. The dataset contains personal events from geohash check-ins and relational events from friendship connections, totaling 8,342,943 events from 196,591 users.
@@
pres-amazon-clothing. The dataset contains Amazon product reviews and ratings in the Clothing, Shoes and Jewelry category, spanning from May 1996 to July 2014. The raw data was originally collected by mcauley2015image. In this dataset, we define personal events as sequences of product IDs and ratings reviewed by a user, for example: B [^0] MLDCZ2:5 and B001OE3F08:3. Relational events represent co-review patterns, where two users have co-reviewed at least three products. The dataset contains event sequences from 185,986 users, with a total of 1,591,947 events.
@@
pres-amazon-electronics. The dataset contains Amazon product reviews and ratings in the Electronics category, originally collected by mcauley2015image. As in pres-amazon-clothing, personal events are defined as sequences of product IDs and ratings, while relational events capture co-review patterns. In total, the dataset contains 2,938,178 events from 254,064 users.
@@
pres-github. This dataset contains GitHub user activity from January 2025, collected from the GH Archive. Personal events include actions such as Push, CreateBranch, CreateRepository, PullRequestOpened, IssuesOpened, and Fork. Relational events represent project collaboration, where two users are linked if both contributed at least five commits or pull requests to the same repository. The dataset includes 102,878,895 events from 3,669,079 users. In this dataset, only personal events include timestamp information; relational events do not have associated timestamps, similar to pres-brightkite and pres-gowalla.
Variability of the datasets.
@@
As shown in Table 1, the PRES datasets vary significantly across multiple aspects. The number of users ranges from around 58 thousand in pres-brightkite to more than 3.5 million in pres-github. The number of events also varies, from approximately 1.5 million in pres-brightkite to over 100 million in pres-github. The ratio between relational and personal events ranges from around 1:3 in pres-gowalla to approximately 1:80 in pres-amazon-clothing. The number of unique events also differs widely, from just 24 in pres-github to more than 1 million in pres-gowalla. In addition, we observe variability in the number of users having personal events, relational events, and both. Some datasets have more users with relational events than with personal events (e.g., pres-brightkite, pres-gowalla), while others show the opposite trend (e.g., pres-amazon-clothing, pres-amazon-electronics, pres-github). These differences in dataset properties present distinct challenges for modeling user events in each dataset.